# Anisotropy of Nanohertz Gravitational Wave Background and Individual Sources from Supermassive Binary Black Holes Based on Cosmological Simulation

### v2

@@ 6 Conclusions and Discussion
It is also interesting to consider that the standard Hellings-Downs curve is derived on the assumption of an isotropic GWB (Hellings & Downs, 1983), the presence of anisotropic components in the GWB could affect the angular correlation, causing it to deviate from the standard curve. It is possible that, based on our results, we can provide a description on HD curve’s configuration under the influence of anisotropy. We leave this also to future work.
@@
We thank Prof. Youjun Lu, Qi Guo, Chengliang Wei for helpful suggestions. Xiao Guo is supported by the Postdoctoral Fellowship Program and of China Postdoctoral Science Foundation under Grant Number BX20230104. We acknowledge the Beijing Super Cloud Computing Center for providing HPC resources that have contributed to the research results reported within this paper (URL: http://www.blsc.cn/).

### v3

@@ 5 Conclusions and Discussion
It is also interesting to consider that the standard Hellings-Downs curve is derived on the assumption of an isotropic GWB (Hellings & Downs, 1983), the presence of anisotropic components in the GWB could affect the angular correlation, causing it to deviate from the standard curve. This has been done in Allen et al. (2024); Grimm et al. (2025) with theoretical framework. It is possible that, based on our results, we can provide a numerical prediction on HD curve’s configuration under the influence of anisotropy. We leave this also to future work.
Appendix A Mass evolution
For computational efficiency, we didn’t consider black hole mass evolution after the merger of galaxies in the main text. In reality, masses of the SMBBHs may continue to grow within the hardening timescale. This appendix quantifies the resulting biases introduced by this simplification. Given that longer hardening timescales permit greater mass growth, we focus solely on the maximum hardening timescale $\tau_{H}=10\,\mathrm{Gyr}$ case to represent the potential upper limit of the impact of mass evolution on our results. We evolve the SMBBHs from the merger of their progenitor galaxies to the GW driven phase with the mass growth rate (including both radio mode and quasar mode) given by the SMBH growth model adopted by SAM (Guo et al. (2011), see also Croton et al. (2006)). We exclude direct mass growth from subsequent black hole mergers to avoid double counting, since as noted in the main text, such events are treated as independent SMBBH sources. For simplicity, we distribute any added mass between the two black holes of a binary pair according to their original mass ratio, thus preserving the mass ratio of the binary system. By tracking the mass evolution process, we derive updated mass parameters for all the SMBBHs. These masses can then be used to compute the characteristic amplitudes, anisotropy and individual source statistics for comparison with results presented in the main text. The results are presented in Table A1 below.
Table A1: Comparison of GWB characteristic amplitude, anisotropy, and resolvable source counts with versus without black hole mass evolution. Here, ME denotes the methodology incorporating black hole mass evolution after galaxy mergers, while NE refers to the approach presented in the main text that excludes such evolutionary effects, and $\rho_{\rm tot}$ is the total energy density, while $\rho_{f_{\rm min}}$ is the energy density in the lowest frequency bin starting from $f_{\rm min}$.
Characteristic amplitude Anisotropy (16%-84% distribution range) Individual source
Models $h_{c}/\rm 0.1yr^{-1}$ $h_{c}/\rm yr^{-1}$ $C_{1}/C_{0}$ of $\rho_{\rm tot}$ $C_{1}/C_{0}$ of $\rho_{\rm f_{\rm min}}$ CPTA SKA-PTA
ME $7.04\times 10^{-16}$ $4.61\times 10^{-18}$ $[7.2\times 10^{-2},6.5\times 10^{-1}]$ $[1.9\times 10^{-2},2.7\times 10^{-1}]$ $1_{-1}^{+1}$ $8_{-4}^{+7}$
NE $3.13\times 10^{-16}$ $2.19\times 10^{-18}$ $[7.7\times 10^{-2},6.2\times 10^{-1}]$ $[2.1\times 10^{-2},2.9\times 10^{-1}]$ $0_{-0}^{+1}$ $2_{-2}^{+3}$
As shown in the table, the characteristic amplitudes increase when considering black hole mass evolution after galaxy mergers, yet they still remain at the same order of magnitude as the previous results. The number of individual sources also shows a slight increase, with a more significant rise seen in the number detectable by SKA-PTA. The variation in the distribution range (16th to 84th percentiles) of anisotropic power spectrum coefficients may stem from multiple factors. On one hand, mass enhancement could amplify the amplitudes of comparatively luminous individual sources, potentially increasing anisotropy. Conversely, growth opportunities for lower-mass systems may narrow mass disparities among sources, potentially reducing anisotropy. The table results demonstrate a broadening of the distribution range for total-energy anisotropy, and a minor overall reduction in anisotropy of energy density within the lowest-frequency bin. The results of both cases are, however, still generally within the same range as those previously reported. Variations in the results of the $\tau_{H}=5\,\mathrm{Gyr}$ case can be inferred from results shown above. Incorporates mass evolution will certainly shift both the characteristic amplitude and resolvable source counts toward values more comparable to the results in the $\tau_{H}=0.1\,\mathrm{Gyr}$ case. Given the shorter hardening timescale relative to the 10 Gyr case, the overall magnitude of variation should be smaller than that shown above. It is interesting to note that, the disparity in GWB properties—characteristic amplitude, anisotropy, and resolvable source counts—between the $\tau_{H}=0.1\,\mathrm{Gyr}$ and $\tau_{H}=5$Gyr cases are significantly smaller than their order-of-magnitude difference in hardening timescales.
@@ Appendix A Mass evolution
We thank Prof. Youjun Lu, Qi Guo, Chengliang Wei for helpful suggestions. Xiao Guo is supported by the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation under Grant Number BX20230104. We acknowledge the Beijing Super Cloud Computing Center for providing HPC resources that have contributed to the research results reported within this paper (URL: http://www.blsc.cn/).